OpenStudy (anonymous):

Find an equation of the line passing through the point (−3, 1) and through the point of intersection of the lines 2x − y + 1 = 0 and 3x + y − 16 = 0. Write your answer in the form y = mx + b.

OpenStudy (raden):

first, u have to find the solution of systems 2x − y + 1 = 0 and 3x + y − 16 = 0. use elimination or substitution method

OpenStudy (anonymous):

i got x=3 And y=7

OpenStudy (raden):

yes... now determine a straight line equation which passes through at points (3,7) and (-3,1) use the formula : (y-y1)/(y2-y1) = (x-x1)/(x2-x1)

OpenStudy (anonymous):

i did all that and got y=x+4 is that my answer

OpenStudy (raden):

yeah, i got it like yours

OpenStudy (anonymous):

ok that's great, thank you so much

OpenStudy (raden):

very welcome